How many ounces of a 25% alcohol solution must be mixed with 18 ounces of a 30% alcohol solution to make a 26% alcohol solution ?
Answer by rothauserc(4718)

let x be the number of ounces of the 25% solution
:
25% solution x 0.25 0.25x
30% solution 18 0.30 5.40
mixture x+18 0.26 0.26(x+18)
:
0.25x + 5.40 = 0.26 * (x + 18)
:
0.25x + 5.40 = 0.26x + 4.68
:
0.01x = 0.72
:
x = 72
:
*************************************************
72 ounces of the 25% solution need to be mixed
*************************************************
: